Dundar Mifflin makes electronic products for the RCMP. The following data is for the first six months:

           

Direct Labor Hours

Manufacturing Overhead

January

45,000

$295,000

February

59,600

$305,575

March

57,000

$323,000

April

58,000

$320,000

May

27,700

$158,400

June

29,000

$152,500

Required:

  1. Use the high-low method to estimate the cost formula                               (3 marks)

  1. Estimate the total overhead cost at an activity level of 43,000 machine hours, using the separate estimates you obtained for its components.                                                    (2 marks)
Answer #1

Variable cost per dlh = (Cost at highest activity - Cost at lowest activity)/(Highest activity-Lowest activity)

= (305,575-158,400)/(59,600-27,700) = 147,175/31,900

= 4.61 per direct labor hour

Fixed cost = Total cost - Total Variable cost

= 305,575 - (59,600*4.61)

= 305,575 - 274,756

= 30,819

Estimated overhead cost for 43,000 machine hours

= Variable (43,000*4.61) + Fixed 30,819

= $229,049